OpenStudy (anonymous):

How do I prove if y=x^n then y^1=n*x^n-1 using lim h-> 0 f(x+h)-f(x) / h

OpenStudy (aravindg):

Hi there! Welcome to Openstudy :) If you have any doubts on how the site works ,feel free to ask me or any other Ambassador with purple A near them .Have great time here . Check out our COC : http://openstudy.com/code-of-conduct

terenzreignz (terenzreignz):

With magic, of course :) LOL Let's see... let \[\large f(x)=x^n \ \ \ n\in \mathbb{Z}^+\]To get its derivative... \[\large f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

So how would we prove it using the definiton of the first derivative?